OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(6, –1) with slope 4. A. y + 6 = 4(x – 1) B. y + 1 = 4(x – 6) C. y = 4x – 1 D. y = 4x – 25 help me someone :/

OpenStudy (anonymous):

Use slope-intercept form, \(y - y_1 = m \left(x-x_1\right)\).

OpenStudy (anonymous):

\(y-(-1) = 4(x - 6)\)

OpenStudy (anonymous):

\(y = 4x - 25\)

OpenStudy (vshiroky):

y+1=4(x-6) y+1=4x-24 y=4x-25

OpenStudy (anonymous):

so the answer is D? i put b or c and my calculations and answer was of course wrong :/ thank you both so much!

OpenStudy (anonymous):

Yes, D.

OpenStudy (anonymous):

thanks! :]